When heated, KClO3 solid forms solid KCl and O2 gas. A sample of KClO3 is heated and 185 mL of gas with a pressure of 775 mmHg is collected over water, at 26 ∘C. At 26 ∘C, the vapor pressure of water is 25 mmHg: 2(KClO3)(s)→2(KCl)(s)+3(O2)(g)
1
Expert's answer
2020-04-28T12:20:04-0400

p1 = 185 mL

V1 = 775 mmHg

p2 = 25 mmHg

V2 -?

Solution:

2KClO3(s)→2KCl(s)+3O2(g)

 At a constant temperature and constant values of the gas mass and its molar mass, the product of the gas volume and its pressure remains constant: pV = const

p1 *V1 = p2*V2

V2 = p1*V1/p2

V2 = 775 mmHg* 185 mL/25 mmHg = 5735 mL =5.735 L

Answer: 5.735 L